### v1

@@ 3.2.4 Dataset statistics
A breakdown of aspect terms by part of speech tags revealed that most were nouns. 45% of all aspect terms were singular nouns such as “fire”, 16% were singular proper nouns such as “London”, and 13% were plural nouns such as “days”. Aside from nouns, 6% of aspect term words were pronouns such as “you”, and 5% were adjectives such as “relevant” (cf. Figure 7). Compared to the ABSA datasets by Mitchell et al. [35] and Dong et al. [36], this new dataset was both less domain-bound and had more diverse part of speech tags. While all emotion classes shared some aspect terms (e.g., “you”, “it”, “fire”), there were words unique to each emotion class, such as "shit" for Anger, "behavior" for Fear, "end" for Sadness, "today" for Happiness, and "month" for None. In the ABEA dataset, aspect terms ranged from 2 to 58 characters, with an average of 10. In comparison, aspect terms in the dataset by Mitchell et al. [35] ranged from 3 to 84 characters, averaging 11.
@@ 3.3.1 GRACE: Gradient-Harmonized and Cascaded Labeling for Aspect-Based Sentiment Analysis
The applicable models are shown in Table 3 and were applied to the Twitter dataset by Mitchell et al. [35]. Most models reported considerably higher performance for the individual sub-tasks compared to compound ATE and ASC. For instance, the GRACE model [2] achieved 75.7% on ATE, while SpanABSA [67] achieved 75.3% and 75.2% for ATE and ASC, respectively. GRACE performed best on Twitter data, with a reported F1-score of 58.3% for joint ATE and ASC and was therefore selected as starting point for ABEA fine-tuning.
@@
[67] SpanABSA BERT-based span-extraction and polarity classification 57.7
[68] Unified-LSTM Unified stacked LSTM with boundary detection and gating 48.0
@@
[2] GRACE Gradient harmonized, cascaded labeling BERT-based model 58.3
[44] MTMVN Multitask Multiview network with CNNs and BiGRUs 47.9
@@
[69] CasNSA Cascade BERT-based model with hierarchical structure 50.1
@@
[70] BDEP-IAM Block-level dependency syntax parsing with interactive attention 58.0
@@
The basic structure of GRACE, which is adapted from the 12-layer, BERT-base model [71], consists of two computational branches for ATE and ASC respectively, which share several root layers. Luo et al. [2] split Bidirectional Encoder Representations from Transformers (BERT) into two branches, sharing the bottom $l$ (9) layers while reserving the top  $l:L$ (3) layers for task-specific ATE and ASC learning. This architecture enables a shared foundational understanding across both branches. Instead of treating ASC as a classification task, both tasks are implemented as sequence labeling to help facilitate co-extraction in a single, unified model. The ATE branch labels tokens as $B$ (beginning-), $I$ (inside-), or $O$ (outside of an aspect term). These labels in turn serve as additional input for the ASC branch, where a modified Transformer decoder generates sentiment labels $POS$ (positive), $NEU$ (neutral), $NEG$ (negative), or $O$ (other words). The passing of ATE label outputs as ASC inputs forms a cascaded labeling approach, which Luo et al. [2] argue enhances the model’s sentiment judgment.
@@
In addition to cascaded labeling, Luo et al. [2] enhance GRACE’s training with the use of Gradient Harmonized Loss Function (GHL) and Virtual Adversarial Training (VAT), both of which are designed to improve model robustness when given difficult or imbalanced training data. GHL, introduced by Li et al. [72], mitigates class imbalance by weighting training examples based on their difficulty according to the model’s predictions during training. This increases model focus on underrepresented, difficult aspect terms, while also preventing overprediction of easier and more frequent neutral tokens. VAT, further enhances model robustness by introducing small adversarial perturbations to the input embeddings during the training process, improving model resilience to data variations [2].
@@
Using these customizations, the fine-tuning was conducted across three training steps, each with a specific number of epochs, learning rate, and setting for VAT and GHL. The first two training steps focused exclusively on ATE. The key difference was that VAT was applied in the second step, but not the first. The third step then incorporated the training of the ASC branch for the full co-extraction of ATE and ASC. Luo et al. [2] configured the hyperparameters based on the SemEval laptop and restaurant datasets. With the extensive customizations for both the model architecture and training process by Luo et al. [2], the GRACE model appeared well suited for the complexities of joint ATE and AEC tasks, learning both tasks in tandem and optimizing the training for robustness.
@@ 3.3.2 Hyperparameter optimization
The GRACE model was initialized in the same manner as described by Luo et al. [2]. Due to the increased classification complexity of emotion classes and considerable structural differences between the SemEval and the new social media ABEA dataset, extensive Hyperparameter Optimization (HPO) was required to identify optimal hyperparameter configurations for ABEA.
The ABEA dataset was split into train, validation, and test subsets, using the same 70%-10%-20% proportions as Luo et al. [2]. In addition to standard hyperparameter categories such as batch size, learning rate, and warm-up proportion, several architectural modifications were tested, including adapting the number of shared layers $l$ for the ATE and ASC branches, the number of decoder blocks in the ASC branch, the classification head architecture, and shallow layer freezing during fine-tuning. Across the three overall training steps, the hyperparameters for the learning rate, GHL-use, VAT-use and batch size were tested individually for each training step. Using the reported optimal ABSA hyperparameter configurations as a baseline [2], hyperparameter configurations were adjusted in both directions where possible. Table 5 gives an overview over selected model specifications.
The adjustment of architectural elements and training settings enable more influence over the model’s fundamental training and structure. First, GHL and VAT configurations were adjusted to evaluate their respective impact on performance. Second, the number of shared layers between the ATE and ASC branches were adjusted in both directions to test the impact of more and less shared learning. When fewer layers are shared, each task-related branch of the GRACE model learns more task-specific representations. Third, the amount of ASC fine-tuning was extended beyond the originally designated epochs of the third training step. Towards this end, fine-tuning iterations were conducted in which ASC was trained from the outset, from the second training step, and as an added fourth training step. In the last approach, all other model layers related to ATE were frozen to conduct independent emotion learning. Fourth, shallow layer freezing was tested by keeping the bottom three, five, and nine BERT layers of the GRACE model frozen for the entirety of the fine-tuning process to preserve the originally learned parameters from pre-training. Lastly, the model’s single linear classification layer was replaced with more complex classifiers. These included a three-layered multilayer perceptron with rectified linear unit activation functions and a 0.1 dropout rate, as well as three, four, and five layered dense classifiers inspired by the DenseNet structure [73], which use linear layers and concatenation of intermediate results to preserve input information while improving feature representation.
@@
Using these metrics, the model’s performance was systematically assessed for each HPO iteration. A baseline comparison was created with the originally reported hyperparameters by Luo et al. [2]. For each hyperparameter category $H_{c}$ where $c\in~{}${$batch~{}sizes,learning~{}rates,...$}, fine-tuning was performed for each configuration $H_{C_{i}}$ where $i\in$$config_{1},config_{2},...config_{n}$}. The averaged performance $P(H_{C_{i}})$ for each configuration was computed from seven sets of macro F1-scores (Equation 1). Specifically, the F1-score was calculated for both ATE and joint ATE +AEC using the validation data $\mathbb{D}_{\text{val}}$ and after each of the three training steps using the test data $\mathbb{D}_{\text{test}}$ (Equation 2). After fine-tuning, additional F1-scores were calculated for ATE utilizing the external ABSA datasets SemEval Laptop $\mathbb{D}_{\text{L}}$ and SemEval Restaurant 2014-16 $\mathbb{D}_{\text{R}}$ [31, 32, 33]. For AEC, a similar validation was performed using the SemEval Affect 2018 dataset $\mathbb{D}_{\text{A}}$ [74]. The performance of each configuration was then assessed against the previous best result. If any configuration from the current hyperparameter category produced a higher average performance than the previously recorded best, that configuration was retained for all subsequent iterations of HPO. Otherwise, the configuration was discarded to maintain the original settings across subsequent iterations.
@@
For validating the AEC task, the SemEval Affect 2018 dataset [74] was used, which consists of sentence-level emotion labels (joy, anger, sadness, fear), as no alternative ABEA validation datasets exist. When comparing the model’s aspect-level emotion inferences with these sentence-level labels, aspect-level emotions were aggregated at the sentence-level, excluding cases with multiple detected emotions per sentence to ensure consistency. To evaluate the co-extraction of ATE and AEC, precision, recall, and the F1-score were determined by focusing only on correctly identified aspect terms. A prediction was considered correct only if both the aspect and its emotion were accurately predicted.
@@ 4 Results
The configurations and results for selected hyperparameters, training, and model adjustments are shown in Table 5. It provides an overview of the model performance using F1-scores, for the new ABEA training dataset and for the validation tests performed on the $\mathbb{D}_{L}$, $\mathbb{D}_{R}$, and $\mathbb{D}_{A}$ datasets. The baseline (configuration 1) is given by the hyperparameters reported by Luo et al. [2] as the optimal settings for ABSA.
@@
Table 6: 10-Fold Cross Validation for Configuration 1 (Original Settings by [2]) and 41 (Highest Performance). The configuration with the highest performance is shown in bold.
@@ 5.1.1 Emotion ambiguity
The inherent subjectivity of human emotions, combined with the lack of contextual information such as facial expressions, spoken tone, or images, makes aspect-level labeling of emotions in written text a considerable challenge. Throughout the annotation process, many Tweets were encountered that could reasonably be labeled with two or more emotions, depending on the Tweet’s unknown context, to which none of the annotators had access. Despite the clear and logically sound structure of the emotion graph proposed by Shaver et al. [1], annotators repeatedly struggled to make definitive decisions on the emotion labels when encountering instances of sarcasm, rhetorical questions, or hypothetical statements. In many cases, sarcasm is self-evident to humans when provided with visual and auditory context. Even in writing, sarcasm can often be detected based through exaggerated emotional expression. Aside from such clear-cut cases, some Tweets fell into the fuzzy boundaries between earnest emotion and sarcasm, posing a challenge for the annotators to confidently label both the emotion and the aspect term. Similarly challenging were hypothetical scenarios, such as “if the Lakers don’t win this game, I might smash something fr”. These types of statements often involved conditional structures that juxtaposed different possible outcomes, each tied to a different emotion. Among other persistent challenges in distinguishing emotions, distinguishing between caring (Happiness) and worry (Fear), or frustration (Anger) and disappointment (Sadness) was frequently debated as conceptually overlapping emotions. The group annotation system and majority vote evaluation helped establish a consensus on these ambiguous cases.
@@ 5.1.4 Inter-annotator agreement
To achieve high IAA, the annotation process incorporated several strategies to enhance consistency among annotators. The annotators attended in-person sessions [10, 40], enabling discussions to resolve ambiguities as they progressed through the dataset. The effectiveness of this annotation approach was reflected by the fact that 89% of all aspect term annotations were included in the final dataset, based on the criteria of being labeled by at least two out of three annotators. This statistic was calculated at the character-level, i.e., the 11% of excluded characters also contained individually annotated whitespaces around aspect term sequences, or additional words adjacent to the final aspect terms. For the emotion labels, each identified aspect term sequence was similarly evaluated for a dominant emotion, which was identified in all but 12 cases that were individually reviewed. This high IAA, which enabled the majority of the initial dataset to be included in the final ABEA training dataset, underscores the value of annotator discussions.
From a social perspective, however, this methodological choice requires critical examination. Group dynamics, especially if pre-existing relationships or authority figures are involved, can bias annotators’ objectivity. Peer conformity theory [75, 76] suggests that individuals may align with dominant voices in group settings to avoid conflict, which can result in “groupthink”, biased labels, or loss of diverse interpretations. To counteract this issue, annotators were repeatedly assured that their own interpretations were valued and that no single person’s opinion was inherently more correct than any other. Any residual effects of group dynamics cannot be quantified precisely and must be accepted as a baseline risk in collaborative annotation.

### v2

@@ 3.2.4 Dataset statistics
A breakdown of aspect terms by part of speech tags revealed that most were nouns. 45% of all aspect terms were singular nouns such as “fire”, 16% were singular proper nouns such as “London”, and 13% were plural nouns such as “days”. Aside from nouns, 6% of aspect term words were pronouns such as “you”, and 5% were adjectives such as “relevant” (cf. Figure 7). Compared to the ABSA datasets by Mitchell.2013 and Dong.2014, this new dataset was both less domain-bound and had more diverse part of speech tags. While all emotion classes shared some aspect terms (e.g., “you”, “it”, “fire”), there were words unique to each emotion class, such as "shit" for Anger, "behavior" for Fear, "end" for Sadness, "today" for Happiness, and "month" for None. In the ABEA dataset, aspect terms ranged from 2 to 58 characters, with an average of 10. In comparison, aspect terms in the dataset by Mitchell.2013 ranged from 3 to 84 characters, averaging 11.
@@ 3.3.1 GRACE: Gradient-Harmonized and Cascaded Labeling for Aspect-Based Sentiment Analysis
The applicable models are shown in Table 3 and were applied to the Twitter dataset by Mitchell.2013. Most models reported considerably higher performance for the individual sub-tasks compared to compound ATE and ASC. For instance, the GRACE model [Luo.2020] achieved 75.7% on ATE, while SpanABSA [Hu.2019] achieved 75.3% and 75.2% for ATE and ASC, respectively. GRACE performed best on Twitter data, with a reported F1-score of 58.3% for joint ATE and ASC and was therefore selected as starting point for ABEA fine-tuning.
@@
[Hu.2019] SpanABSA BERT-based span-extraction and polarity classification 57.7
[Li.2019a] Unified-LSTM Unified stacked LSTM with boundary detection and gating 48.0
@@
[Luo.2020] GRACE Gradient harmonized, cascaded labeling BERT-based model 58.3
[Bie.2021] MTMVN Multitask Multiview network with CNNs and BiGRUs 47.9
@@
[Ding.2022] CasNSA Cascade BERT-based model with hierarchical structure 50.1
@@
[Xiang.2023] BDEP-IAM Block-level dependency syntax parsing with interactive attention 58.0
@@
The basic structure of GRACE, which is adapted from the 12-layer, BERT-base model [Devlin.2019], consists of two computational branches for ATE and ASC respectively, which share several root layers. Luo.2020 split Bidirectional Encoder Representations from Transformers (BERT) into two branches, sharing the bottom $l$ (9) layers while reserving the top  $l:L$ (3) layers for task-specific ATE and ASC learning. This architecture enables a shared foundational understanding across both branches. Instead of treating ASC as a classification task, both tasks are implemented as sequence labeling to help facilitate co-extraction in a single, unified model. The ATE branch labels tokens as $B$ (beginning-), $I$ (inside-), or $O$ (outside of an aspect term). These labels in turn serve as additional input for the ASC branch, where a modified Transformer decoder generates sentiment labels $POS$ (positive), $NEU$ (neutral), $NEG$ (negative), or $O$ (other words). The passing of ATE label outputs as ASC inputs forms a cascaded labeling approach, which Luo.2020 argue enhances the model’s sentiment judgment.
@@
In addition to cascaded labeling, Luo.2020 enhance GRACE’s training with the use of Gradient Harmonized Loss Function (GHL) and Virtual Adversarial Training (VAT), both of which are designed to improve model robustness when given difficult or imbalanced training data. GHL, introduced by Li.2018a, mitigates class imbalance by weighting training examples based on their difficulty according to the model’s predictions during training. This increases model focus on underrepresented, difficult aspect terms, while also preventing overprediction of easier and more frequent neutral tokens. VAT, further enhances model robustness by introducing small adversarial perturbations to the input embeddings during the training process, improving model resilience to data variations [Luo.2020].
@@
Using these customizations, the fine-tuning was conducted across three training steps, each with a specific number of epochs, learning rate, and setting for VAT and GHL. The first two training steps focused exclusively on ATE. The key difference was that VAT was applied in the second step, but not the first. The third step then incorporated the training of the ASC branch for the full co-extraction of ATE and ASC. Luo.2020 configured the hyperparameters based on the SemEval laptop and restaurant datasets. With the extensive customizations for both the model architecture and training process by Luo.2020, the GRACE model appeared well suited for the complexities of joint ATE and AEC tasks, learning both tasks in tandem and optimizing the training for robustness.
@@ 3.3.2 Hyperparameter optimization
The GRACE model was initialized in the same manner as described by Luo.2020. Due to the increased classification complexity of emotion classes and considerable structural differences between the SemEval and the new social media ABEA dataset, extensive Hyperparameter Optimization (HPO) was required to identify optimal hyperparameter configurations for ABEA.
The ABEA dataset was split into train, validation, and test subsets, using the same 70%-10%-20% proportions as Luo.2020. In addition to standard hyperparameter categories such as batch size, learning rate, and warm-up proportion, several architectural modifications were tested, including adapting the number of shared layers $l$ for the ATE and ASC branches, the number of decoder blocks in the ASC branch, the classification head architecture, and shallow layer freezing during fine-tuning. Across the three overall training steps, the hyperparameters for the learning rate, GHL-use, VAT-use and batch size were tested individually for each training step. Using the reported optimal ABSA hyperparameter configurations as a baseline [Luo.2020], hyperparameter configurations were adjusted in both directions where possible. Table 5 gives an overview over selected model specifications.
The adjustment of architectural elements and training settings enable more influence over the model’s fundamental training and structure. First, GHL and VAT configurations were adjusted to evaluate their respective impact on performance. Second, the number of shared layers between the ATE and ASC branches were adjusted in both directions to test the impact of more and less shared learning. When fewer layers are shared, each task-related branch of the GRACE model learns more task-specific representations. Third, the amount of ASC fine-tuning was extended beyond the originally designated epochs of the third training step. Towards this end, fine-tuning iterations were conducted in which ASC was trained from the outset, from the second training step, and as an added fourth training step. In the last approach, all other model layers related to ATE were frozen to conduct independent emotion learning. Fourth, shallow layer freezing was tested by keeping the bottom three, five, and nine BERT layers of the GRACE model frozen for the entirety of the fine-tuning process to preserve the originally learned parameters from pre-training. Lastly, the model’s single linear classification layer was replaced with more complex classifiers. These included a three-layered multilayer perceptron with rectified linear unit activation functions and a 0.1 dropout rate, as well as three, four, and five layered dense classifiers inspired by the DenseNet structure [Huang.2018a], which use linear layers and concatenation of intermediate results to preserve input information while improving feature representation.
@@
Using these metrics, the model’s performance was systematically assessed for each HPO iteration. A baseline comparison was created with the originally reported hyperparameters by Luo.2020. For each hyperparameter category $H_{c}$ where $c\in$~{batchsizes,learningrates,batch~sizes,learning~rates,...}, fine-tuning was performed for each configuration $H_{C_{i}}$ where $i\in$ $config_{1},config_{2},...config_{n}$}. The averaged performance $P(H_{C_{i}})$ for each configuration was computed from seven sets of macro F1-scores (Equation 1). Specifically, the F1-score was calculated for both ATE and joint ATE +AEC using the validation data $\mathbb{D}_{\text{val}}$ and after each of the three training steps using the test data $\mathbb{D}_{\text{test}}$ (Equation 2). After fine-tuning, additional F1-scores were calculated for ATE utilizing the external ABSA datasets SemEval Laptop $\mathbb{D}_{\text{L}}$ and SemEval Restaurant 2014-16 $\mathbb{D}_{\text{R}}$ [Pontiki.2014, Pontiki.2015, Pontiki.2016a]. For AEC, a similar validation was performed using the SemEval Affect 2018 dataset $\mathbb{D}_{\text{A}}$ [Mohammad.2018]. The performance of each configuration was then assessed against the previous best result. If any configuration from the current hyperparameter category produced a higher average performance than the previously recorded best, that configuration was retained for all subsequent iterations of HPO. Otherwise, the configuration was discarded to maintain the original settings across subsequent iterations.
@@
For validating the AEC task, the SemEval Affect 2018 dataset [Mohammad.2018] was used, which consists of sentence-level emotion labels (joy, anger, sadness, fear), as no alternative ABEA validation datasets exist. When comparing the model’s aspect-level emotion inferences with these sentence-level labels, aspect-level emotions were aggregated at the sentence-level, excluding cases with multiple detected emotions per sentence to ensure consistency. To evaluate the co-extraction of ATE and AEC, precision, recall, and the F1-score were determined by focusing only on correctly identified aspect terms. A prediction was considered correct only if both the aspect and its emotion were accurately predicted.
@@ 4 Results
The configurations and results for selected hyperparameters, training, and model adjustments are shown in Table 5. It provides an overview of the model performance using F1-scores, for the new ABEA training dataset and for the validation tests performed on the $\mathbb{D}_{L}$, $\mathbb{D}_{R}$, and $\mathbb{D}_{A}$ datasets. The baseline (configuration 1) is given by the hyperparameters reported by Luo.2020 as the optimal settings for ABSA.
@@
Table 6: 10-Fold Cross Validation for Configuration 1 (Original Settings by [Luo.2020]) and 41 (Highest Performance). The configuration with the highest performance is shown in bold.
@@ 5.1.1 Emotion ambiguity
The inherent subjectivity of human emotions, combined with the lack of contextual information such as facial expressions, spoken tone, or images, makes aspect-level labeling of emotions in written text a considerable challenge. Throughout the annotation process, many Tweets were encountered that could reasonably be labeled with two or more emotions, depending on the Tweet’s unknown context, to which none of the annotators had access. Despite the clear and logically sound structure of the emotion graph proposed by Shaver.1987, annotators repeatedly struggled to make definitive decisions on the emotion labels when encountering instances of sarcasm, rhetorical questions, or hypothetical statements. In many cases, sarcasm is self-evident to humans when provided with visual and auditory context. Even in writing, sarcasm can often be detected based through exaggerated emotional expression. Aside from such clear-cut cases, some Tweets fell into the fuzzy boundaries between earnest emotion and sarcasm, posing a challenge for the annotators to confidently label both the emotion and the aspect term. Similarly challenging were hypothetical scenarios, such as “if the Lakers don’t win this game, I might smash something fr”. These types of statements often involved conditional structures that juxtaposed different possible outcomes, each tied to a different emotion. Among other persistent challenges in distinguishing emotions, distinguishing between caring (Happiness) and worry (Fear), or frustration (Anger) and disappointment (Sadness) was frequently debated as conceptually overlapping emotions. The group annotation system and majority vote evaluation helped establish a consensus on these ambiguous cases.
@@ 5.1.4 Inter-annotator agreement
To achieve high IAA, the annotation process incorporated several strategies to enhance consistency among annotators. The annotators attended in-person sessions [Luo.2022, Maitama.2021], enabling discussions to resolve ambiguities as they progressed through the dataset. The effectiveness of this annotation approach was reflected by the fact that 89% of all aspect term annotations were included in the final dataset, based on the criteria of being labeled by at least two out of three annotators. This statistic was calculated at the character-level, i.e., the 11% of excluded characters also contained individually annotated whitespaces around aspect term sequences, or additional words adjacent to the final aspect terms. For the emotion labels, each identified aspect term sequence was similarly evaluated for a dominant emotion, which was identified in all but 12 cases that were individually reviewed. This high IAA, which enabled the majority of the initial dataset to be included in the final ABEA training dataset, underscores the value of annotator discussions.
From a social perspective, however, this methodological choice requires critical examination. Group dynamics, especially if pre-existing relationships or authority figures are involved, can bias annotators’ objectivity. Peer conformity theory [Asch.1951, Asch.1955] suggests that individuals may align with dominant voices in group settings to avoid conflict, which can result in “groupthink”, biased labels, or loss of diverse interpretations. To counteract this issue, annotators were repeatedly assured that their own interpretations were valued and that no single person’s opinion was inherently more correct than any other. Any residual effects of group dynamics cannot be quantified precisely and must be accepted as a baseline risk in collaborative annotation.